Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis of standard direct labor-hours. The budgeted variable manufacturing overhead is $4.40 per direct labor-hour and the budgeted fixed manufacturing overhead is $1,764,000 per year. 


The standard quantity of materials is 4 pounds per unit and the standard cost is $9.00 per pound. The standard direct labor- hours per unit is 1.5 hours and the standard labor rate is $13.20 per hour. 


The company planned to operate at a denominator activity level of 210,000 direct labor-hours and to produce 140,000 units of product during the most recent year. Actual activity and costs for the year were as follows: 

Actual number of units produced168,000
Actual direct labor-hours worked273,000
Actual variable manufacturing overhead cost incurred$709,800
Actual fixed manufacturing overhead cost incurred$1,911,000

 

Required:

 1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements.

 2. Prepare a standard cost card for the company's product.

 3a. Compute the standard direct labor-hours allowed for the year's production.

 3b. Complete the following Manufacturing Overhead T-account for the year.

 4. Determine the reason for any underapplied or overapplied overhead for the year by computing the variable overhead rate and efficiency variances and the fixed overhead budget and volume variances.


Complete this question by entering your answers in the tabs below. Req 1 Req 2 Req ЗА Req 3B Req 4 Compute the predeterminedComplete this question by entering your answers in the tabs below. Req 1 Reg 1 Reg 2 Req 3A Req 3A Req 3B Req 4 Prepare a staComplete this question by entering your answers in the tabs below. Reg 1 Req 2 Req 3A Req 3B Reg 4 Compute the standard direcComplete this question by entering your answers in the tabs below. Reg 1 Req 2 Req ЗА Req 3B Req 4 Complete the following ManComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req ЗА Req 3B Req 4 Determine the reason for a
Answer #1

1) Predetermined Overhead rate = 4.40 + (1,764,000/210,000) = 12.8 Per DLH

Variable Rate = 4.40 Per DLH

Fixed Rate = 8.40 Per DLH

2)

Direct Materials           4.00 Pounds at           9.00 Per Pound         36.00
Direct Labor           1.50 DLHs at         13.20 Per DLH         19.80
Variable Overhead           1.50 DLHs at           4.40 Per DLH           6.60
Fixed Overhead           1.50 DLHs at           8.40 Per DLH         12.60
Standard cost per unit         75.00

3a) Standard Direct Labor Hours allowed = 168,000 x 1.5 = 252,000 Hours

3b)

Manufacturing Overhead
Actual Cost    2,620,800.00    3,225,600.00 Applied Cost
       604,800.00 Over-applied overhead

4)Variable Overhead rate Variance = 709,800 - (273,000 x 4.4) = 491,400 (F)

Variable overhead efficiency variance = (273,000 - 252,000) x 4.4 = 92,400 (U)

Fixed Overhead Budget Variance = 1,911,000 - 1,764,000 = 147,000 (U)

Fixed Overhead Volume Variance = (210,000 - 252,000) x 8.4 = 352,800 (F)